# GVE-LPA: Fast Label Propagation Algorithm (LPA) for Community Detection in Shared Memory Setting

### v5

@@ Abstract.
Community detection is the problem of identifying natural divisions in networks. Efficient parallel algorithms for this purpose are crucial in various applications, particularly as datasets grow to substantial scales. This technical report presents an optimized parallel implementation of the Label Propagation Algorithm (LPA), a high speed community detection method, for shared memory multicore systems. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, our LPA, which we term as GVE-LPA, outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × respectively - achieving a processing rate of $1.4B$ edges/s on a $3.8B$ edge graph. In addition, GVE-LPA scales at a rate of 1.7×1.7$\times$1.7 × every doubling of threads.
@@ 1. Introduction
Community detection is the problem of identifying cohesive groups of vertices, also known as clusters, in complex networks that minimize cuts and maximize internal density (Garza and Schaeffer, 2019). This problem finds applications in various domains, including drug discovery, protein annotation, topic exploration, anomaly detection, and criminal identification. Communities identified are intrinsic when based on network topology alone, and are disjoint when each vertex belongs to only one community (Gregory, 2010). One of the difficulties in the community detection problem is the lack of apriori knowledge on the number and size distribution of communities (Blondel et al., 2008).
@@
The Label Propagation Algorithm (LPA), also known as RAK, (Raghavan et al., 2007) is a popular diffusion-based approach for identifying communities. It is faster and more scalable than the Louvain method (Blondel et al., 2008), another high-quality community detection algorithm, as it does not require repeated optimization steps and is easier to parallelize (Newman, 2004; Raghavan et al., 2007). Due to this, LPA has been used in a number of other graph problems and applications, such as, finding connected components (Stergiou et al., 2018), graph partitioning (Slota et al., 2014; Wang et al., 2014; Meyerhenke et al., 2014, 2016, 2017; Bae et al., 2020; Akhremtsev et al., 2020; Slota et al., 2020; Zhang et al., 2020a), hypergraph partitioning (Henne, 2015; Gottesbüren et al., 2021), graph coarsening (Valejo et al., 2020), vertex reordering and graph compression (Boldi et al., 2011), unsupervised part-of-speech tagging (Das and Petrov, 2011), sectionalizing power systems (Aziz et al., 2023). Further, Peng et al. (Peng et al., 2014) observe that, while LPA achieves lower modularity scores (a metric for assessing community quality (Newman, 2004)) than the Louvain method, it attains the highest NMI score when compared to the ground truth. A large number of LPA variants have also been proposed (Li et al., 2015a; Farnadi et al., 2015; Li et al., 2015b; SHEN and YANG, 2016; Mohan et al., 2017; Zhang et al., 2017; Berahmand and Bouyer, 2018a; Ma et al., 2018; Sattari and Zamanifar, 2018b; Zheng et al., 2018; Xu et al., 2019; Zarei et al., 2020; Maleki et al., 2020; Zhang et al., 2020b; El Kouni et al., 2021; Roghani et al., 2021; Zhang et al., 2023). Existing studies on LPA propose a number of algorithmic optimizations and parallelization techniques, but there continue to be a number of issue with each implementation. Further, the race for AI development has caused the prices of GPUs to skyrocket. These issues, and the large applicability of LPA, motivates us to present GVE-LPA, [^1] https://github.com/puzzlef/rak-communities-openmp an efficient implementation of LPA for multicore architectures.
@@ 2. Related work
A few open source implementations and software packages have been developed for community detection using LPA. Fast Label Propagation Algorithm (FLPA) (Traag and Šubelj, 2023) is a fast variant of the LPA, which utilizes a queue-based approach to process only vertices with recently updated neighborhoods. NetworKit (Staudt et al., 2016) is a software package designed for analyzing the structural aspects of graph data sets with billions of connections. It is implemented as a hybrid with C++ kernels and a Python frontend, and includes parallel implementation of LPA. igraph (Csardi et al., 2006) is a similar package, written in C, with Python, R, and Mathematica frontends. It is widely used in academic research, and includes implementation of LPA.
@@
We now discuss the implementation of LPA in NetworKit, igraph, and FLPA. The parallel implementation of LPA in NetworKit is given in NetworKit::PLP::run(). This function starts with a unique label for each node. To check for convergence, they use a tolerance of $10^{-5}$, i.e., the algorithm converges once the labels of less than $0.001\%$ of vertices change. This is also referred to as threshold heuristic. To track active nodes, a boolean flag vector is employed, serving as a vertex pruning optimization. A parallel for loop, which uses OpenMP’s guided schedule, then processes only the active nodes. For storing label weights, an std::map is used for each vertex. In contarst, the LPA implementation of igraph and FLPA is sequential. igraph’s LPA implementation, found in the function igraph_community_label_propagation(), shuffles the node order in each iteration. It keeps track of the dominant labels for each new neighbor of a vertex and selects one of the dominant labels using a random number generator. The hashtable is cleared by iterating through the neighbors. Unlike NetworKit, igraph does not use the pruning optimization. It alternates between label updating and control iterations, checking if the current label of a node is not dominant during control iterations, with no tolerance for non-dominant labels. The FLPA implementation in igraph, detailed in the function igraph_i_community_label_propagation(), uses a dequeue for managing a set of unprocessed vertices, which indicates convergence when empty. FLPA does not shuffle for random node order. If a label change occurs, FLPA considers neighbors that are not in the same community.
@@
However, we identify a few issues with the LPA implementations of NetworKit, igraph, and FLPA. In order to assign a unique label to each node in the graph, NetworKit uses OpenMP’s plain parallel for, i.e., it uses static loop scheduling with a chunk size of $1$. This can result in false sharing as threads make writes to consecutive locations in memory. Next, to keep track of the weights associated with each label for a given vertex, they use an std::map. Our experiments show that this is quite inefficient. We, in contrast, use a per-thread keys list and a per-thread full-size values array, as our hashtable. In order to check for convergence, NetworKit uses a tolerance of $10^{-5}$. However, we observe that a tolerance of $10^{-2}$ is generally obtains community of nearly the same quality (in terms of modularity), but converges much faster. Further, NetworKit uses atomic operations on a shared variable to keep a count of the number of updated vertices. This can introduce contention. We, instead, make use of parallel reduce for this. Finally, NetworKit uses a boolean flag vector to keep track of keep track of the set of active nodes. However, our experiments show that using an 8-bit integer flag vector is more efficient, despite its larger memory footprint.
@@
We now discuss issues in the LPA implementation of igraph. igraph randomly shuffles the processing order of nodes in each iteration. This can get quite expensive. igraph checks for convergence in alternate iterations, and algorithm is considered to have converged only if the labels of all nodes are dominant, i.e., are the most popular. Thus it can take a large number of iterations to converge, with minimal gain in community quality. In addition, igraph does not use vertex pruning optimization, i.e., it does not maintain the set of active nodes. Finally, given multiple dominant labels, igraph select a random dominant label as the label of the given node — however, random number generation is slow. We instead use the dominant label with the smallest ID as the label of the given vertex.
@@
Finally, we discuss the issues in FLPA’s implementation. Similar to igraph LPA, given multiple dominant labels, FLPA select a random dominant label as the label of the given node — however, random number generation is slow. Further, FLPA considers the algorithm to have converged only when there are no active nodes. Thus it can take a large number of iterations to converge, with minimal gain in community quality.
@@ 4.1. Optimizations for LPA
We use a parallel implementation of LPA and experiment with different optimizations and parameter settings. We use the asynchronous version of LPA, wherein threads operate independently on distinct regions of the graph. This facilitates quicker convergence, but may introduce greater variability in the final result. Further, we observe that parallel LPA obtains communities of higher quality than its sequential implementation, possibly due to randomization. We allocate a separate hashtable per thread to keep track of the total weight of each unique label linked to a vertex.
We now evaluate multiple alternatives for a few optimizations, and show the relative time and the relative modularity of communities obtained by each alternative in Figure 3. We perform these tests on every graph in the dataset (refer to Table 1), conducting them five times on each graph to minimize the influence of noise. We then calculate the geometric mean for the runtime and arithmetic mean for the modularity, and represent them as ratios within each optimization category. While none of our optimizations are particularly novel, we carefully address each one and combine them to achieve superior performance compared to other implementations of LPA. We believe our implementation will aid the community in advancing more rapidly and tackling more challenges.
@@ 4.1.1. Adjusting OpenMP loop schedule
We attempt static, dynamic, guided, and auto loop scheduling approaches of OpenMP (each with a chunk size of $2048$) to parallelize LPA. We consider OpenMP’s dynamic loop schedule to be the best choice, due to its ability of work balancing among threads, and because it yields a runtime reduction of $27\%$ when compared to OpenMP’s auto loop schedule, while incurring only a $0.7\%$ reduction in the modularity of obtained communities (likely to be just noise).
@@ 5.2. Comparing Performance of GVE-LPA
Figure 4(a) shows the runtimes of FLPA, igraph LPA, NetworKit LPA, and GVE-LPA for each graph in the dataset, while Figure 4(b) shows the speedup of GVE-LPA with respect to each aforementioned implementation. igraph LPA fails to run on protein k-mer graphs, i.e., kmer_A2a and kmer_V1r, and hence its results on these graphs are not shown. GVE-LPA exhibits an average speedup of 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × over FLPA, igraph LPA, and NetworKit LPA respectively. While FLPA and igraph LPA are sequential implementations, GVE-LPA, running on $64$ threads, is more than 64×64$\times$64 × faster than these implementations. On the sk-2005 graph, GVE-LPA identifies communities in $2.7$ seconds, achieving a processing rate of $1.4$ billion edges/s. Figure 4(c) shows the modularity of communities obtained by each implementation. On average, GVE-LPA obtains $6.6\%$ / $0.2\%$ higher modularity than FLPA (especially on road networks) and igraph LPA respectively, and $4.1\%$ lower modularity than NetworKit LPA (especially on protein k-mer graphs).
@@
Next, we proceed to compare the performance of GVE-LPA with GVE-Louvain in Figure 5. Figures 5(a), 5(b), and 5(c) present the runtimes, speedup (of GVE-LPA with respect to GVE-Louvain), and modularity of GVE-Louvain and GVE-LPA for each graph in the dataset. GVE-LPA offers a mean speedup of 5.4×5.4$\times$5.4 × compared to GVE-Louvain, particularly on social networks, road networks, and protein k-mer graphs, while achieving on average $10.9\%$ lower modularity, especially on social networks and protein k-mer graphs.
In Figure 6, we assess the runtime per edge of GVE-LPA. Results indicate that GVE-LPA exhibits a higher $\text{runtime}/|E|$ factor on graphs with a low average degree (road networks and protein k-mer graphs) and on graphs with a poor community structure (com-LiveJournal and com-Orkut).
@@ 6. Conclusion
In summary, this study focused on optimizing the Label Propagation Algorithm (LPA), a high-speed community detection algorithm, in the shared memory setting. Comparative assessments against competitive open-source implementations (FLPA) and packages (igraph and NetworKit) indicate that GVE-LPA outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × respectively, while identifying communities of $6.6\%$ / $0.2\%$ higher quality than FLPA and igraph LPA respectively, and of $4.1\%$ lower quality than NetworKit LPA. On a web graph with $3.8$ billion edges, GVE-LPA identifies communities in $2.7$ seconds, and thus achieves a processing rate of $1.4$ billion edges/s. GVE-LPA is ideal for applications requiring a high-speed clustering algorithm while accepting a compromise in clustering quality, as it is on average 5.4×5.4$\times$5.4 × faster than GVE-Louvain, has a strong scaling factor of 1.7×1.7$\times$1.7 × for every doubling of threads, but identifies communities with on average $10.9\%$ lower modularity than GVE-Louvain.

### v6

@@ Abstract.
Community detection is the problem of identifying natural divisions in networks. Efficient parallel algorithms for this purpose are crucial in various applications, particularly as datasets grow to substantial scales. This technical report presents an optimized parallel implementation of the Label Propagation Algorithm (LPA), a high speed community detection method, for shared memory multicore systems. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, our LPA, which we term as GVE-LPA, outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97000×97000$\times$97000 ×, and 40×40$\times$40 × respectively - achieving a processing rate of $1.4B$ edges/s on a $3.8B$ edge graph. In addition, GVE-LPA scales at a rate of 1.7×1.7$\times$1.7 × every doubling of threads.
@@ 1. Introduction
Community detection is the problem of identifying cohesive groups of vertices, also known as clusters, in complex networks that minimize cuts and maximize internal density (Garza and Schaeffer, 2019). Communities identified are intrinsic when based on network topology alone, and are disjoint when each vertex belongs to only one community (Gregory, 2010). One of the difficulties in community detection is the lack of apriori knowledge on the number and size distribution of communities (Blondel et al., 2008).
Community detection finds applications in various domains. In drug discovery, it helps identify groups of similar compounds or target proteins, thereby facilitating the discovery of new therapeutic agents (Ma et al., 2019; Udrescu et al., 2020). In the health domain, it aids in understanding the dynamics of groups susceptible to epidemic diseases (Salathé and Jones, 2010), detecting diseases like lung cancer (Bechtel et al., 2005), and categorizing tumor types using genomic datasets (Haq and Wang, 2016). Additionally, it has been applied to elucidate the structure and evolution of metabolic networks (Pfeiffer et al., 2005; Kim et al., 2009), Gene Regulatory Networks (GRNs) (Rivera et al., 2010), and Lateral Gene Transfer (LGT) networks (Popa et al., 2011). Another significant application of community detection is in the analysis of human brain networks (Bullmore and Sporns, 2009; He and Evans, 2010). In ecological studies, it is used to determine if food webs are organized into compartments, where species within the same compartment frequently interact among themselves but have fewer interactions with species in different compartments (Krause et al., 2003; Rezende et al., 2009; Guimerà et al., 2010).
@@
The Label Propagation Algorithm (LPA), also known as RAK, (Raghavan et al., 2007) is a popular diffusion-based approach for identifying communities. It is faster and more scalable than the Louvain method (Blondel et al., 2008), another high-quality community detection algorithm, as it does not require repeated optimization steps and is easier to parallelize (Newman, 2004; Raghavan et al., 2007). Due to this, LPA has been used in a number of other graph problems and applications, such as, finding connected components (Stergiou et al., 2018), graph partitioning (Slota et al., 2014; Wang et al., 2014; Meyerhenke et al., 2014, 2016, 2017; Bae et al., 2020; Akhremtsev et al., 2020; Slota et al., 2020; Zhang et al., 2020a), hypergraph partitioning (Henne, 2015; Gottesbüren et al., 2021), graph coarsening (Valejo et al., 2020), vertex reordering and graph compression (Boldi et al., 2011), unsupervised part-of-speech tagging (Das and Petrov, 2011), sectionalizing power systems (Aziz et al., 2023). Further, Peng et al. (Peng et al., 2014) observe that, while LPA achieves lower modularity scores (a metric for assessing community quality (Newman, 2004)) than the Louvain method — a lower modularity score for LPA implies that it may struggle to properly cluster nodes that are not well-connected to any particular community— it attains the highest Normalized Mutual Information (NMI) score when compared to the ground truth.
A large number of LPA variants have also been proposed (Li et al., 2015a; Farnadi et al., 2015; Li et al., 2015b; SHEN and YANG, 2016; Mohan et al., 2017; Zhang et al., 2017; Berahmand and Bouyer, 2018a; Ma et al., 2018; Sattari and Zamanifar, 2018b; Zheng et al., 2018; Xu et al., 2019; Zarei et al., 2020; Maleki et al., 2020; Zhang et al., 2020b; El Kouni et al., 2021; Roghani et al., 2021; Zhang et al., 2023). Existing studies on LPA propose a number of algorithmic optimizations and parallelization techniques, but there continue to be a number of issue with each implementation. Further, the race for AI development has caused the prices of GPUs to skyrocket. These issues, and the large applicability of LPA, motivates us to present GVE-LPA, [^1] https://github.com/puzzlef/rak-communities-openmp an efficient implementation of LPA for multicore architectures.
@@ 2. Related work
A few open source implementations have been developed for community detection using LPA. Fast Label Propagation Algorithm (FLPA) (Traag and Šubelj, 2023) is a fast variant of the LPA, which utilizes a queue-based approach to process only vertices with recently updated neighborhoods. NetworKit (Staudt et al., 2016) is a software package designed for analyzing the structural aspects of graph data sets with billions of connections. It is implemented as a hybrid with C++ kernels and a Python frontend, and includes parallel implementation of LPA. igraph (Csardi et al., 2006) is a similar package, written in C, with Python, R, and Mathematica frontends. It is widely used in research, and includes an implementation of LPA.
@@
We now discuss the implementation of LPA in NetworKit, igraph, and FLPA. The parallel implementation of LPA in NetworKit is given in NetworKit::PLP::run(). This function starts with a unique label for each node. To check for convergence, they use a tolerance of $10^{-5}$, i.e., the algorithm converges once the labels of less than $0.001\%$ of vertices change. This is also referred to as threshold heuristic. To track active nodes, a boolean flag vector is employed, serving as a vertex pruning optimization. A parallel for loop, which uses OpenMP’s guided schedule, then processes only the active nodes. For storing label weights, an std::map is used for each vertex. In contrast, the LPA implementation of igraph and FLPA is sequential. igraph’s LPA implementation, found in the function igraph_community_label_propagation(), shuffles the node order in each iteration. It keeps track of the dominant labels for each new neighbor of a vertex and selects one of the dominant labels using a random number generator. The hashtable is cleared by iterating through the neighbors. Unlike NetworKit, igraph does not use the pruning optimization. It alternates between label updating and control iterations, checking if the current label of a node is not dominant during control iterations, with no tolerance for non-dominant labels. The FLPA implementation in igraph, detailed in the function igraph_i_community_label_propagation(), uses a dequeue for managing a set of unprocessed vertices, which indicates convergence when empty. FLPA does not shuffle for random node order. If a label change occurs, FLPA considers neighbors that are not in the same community.
@@
However, we identify a few issues with the LPA implementations of NetworKit, igraph, and FLPA. In order to assign a unique label to each node in the graph, NetworKit uses OpenMP’s plain parallel for, i.e., it uses static loop scheduling with a chunk size of $1$. This can result in false sharing as threads make writes to consecutive locations in memory. Next, to keep track of the weights associated with each label for a given vertex, they use an std::map. Our experiments show that this is quite inefficient. We, in contrast, use a per-thread keys list and a per-thread full-size values array, as our hashtable. In order to check for convergence, NetworKit uses a tolerance of $10^{-5}$. However, we observe that a tolerance of $10^{-2}$ is generally obtains community of nearly the same quality (in terms of modularity), but converges much faster. Further, NetworKit uses atomic operations on a shared variable to keep a count of the number of updated vertices. This can introduce contention. We, instead, make use of parallel reduce for this. Finally, NetworKit uses a boolean vector to keep track of the set of active nodes. However, our experiments show that using an 8-bit integer flag vector is more efficient.
@@
We now discuss issues in the LPA implementation of igraph. igraph randomly shuffles the processing order of nodes in each iteration. This can get quite expensive. igraph checks for convergence in alternate iterations, and algorithm is considered to have converged only if the labels of all nodes are dominant, i.e., are the most popular. Thus it can take a large number of iterations to converge, with minimal gain in community quality. In addition, igraph does not use vertex pruning optimization, i.e., it does not maintain the set of active nodes. Finally, given multiple dominant labels, igraph select a random dominant label as the label of the given node — however, random number generation is slow. We use the dominant label with the smallest ID as the label of the given vertex.
@@
Finally, we discuss the issues in FLPA’s implementation. Similar to igraph LPA, given multiple dominant labels, FLPA selects a random dominant label as the label of the given node — again, random number generation is slow. Further, FLPA considers the algorithm to have converged only when there are no active nodes. Thus it can take a large number of iterations to converge, with minimal gain in community quality.
@@ 4.1. Optimizations for LPA
We use an asynchronous parallel implementation of LPA, wherein threads operate independently on distinct regions of the graph. This facilitates quicker convergence, but may introduce greater variability in the final result. Further, we observe that parallel LPA obtains communities of higher quality than its sequential implementation, possibly due to randomization. We allocate a separate hashtable per thread to keep track of the total weight of each unique label linked to a vertex.
We now evaluate alternatives for a few optimizations, on the system mentioned in Section 5.1, for every graph in the dataset (Table 1) — conducting them five times on each graph to minimize the influence of noise. We then calculate the geometric mean for the runtime and arithmetic mean for the modularity, and represent them as ratios within each optimization category, as shown in Figure 3. While none of our optimizations are particularly novel, we carefully address each one and combine them to achieve superior performance compared to other implementations of LPA.
@@ 4.1.1. Adjusting OpenMP loop schedule
We attempt various loop scheduling approaches of OpenMP — each using a chunk size of $2048$ (we observe this strikes a balance between minimizing load imbalance and reducing scheduling overhead) to parallelize LPA. We consider OpenMP’s dynamic loop schedule to be the best choice, due to its ability of work balancing among threads, and because it yields a runtime reduction of $27\%$ when compared to OpenMP’s auto loop schedule, while incurring only a $0.7\%$ reduction in the modularity of obtained communities (likely to be just noise).
@@ 5.2. Comparing Performance of GVE-LPA
Figure 4(a) shows the runtimes of FLPA, igraph LPA, NetworKit LPA, and GVE-LPA for each graph in the dataset, while Figure 4(b) shows the speedup of GVE-LPA with respect to each aforementioned implementation. igraph LPA fails to run on protein k-mer graphs, i.e., kmer_A2a and kmer_V1r, and hence its results on these graphs are not shown. GVE-LPA exhibits an average speedup of 139×139$\times$139 ×, 97000×97000$\times$97000 ×, and 40×40$\times$40 × over FLPA, igraph LPA, and NetworKit LPA respectively. While FLPA and igraph LPA are sequential implementations, GVE-LPA, running on $64$ threads, is more than 64×64$\times$64 × faster than these implementations. On the sk-2005 graph, GVE-LPA identifies communities in $2.7$ seconds, achieving a processing rate of $1.4$ billion edges/s. Figure 4(c) shows the modularity of communities obtained by each implementation. On average, GVE-LPA obtains $6.6\%$ / $0.2\%$ higher modularity than FLPA (especially on road networks) and igraph LPA respectively, and $4.1\%$ lower modularity than NetworKit LPA (especially on protein k-mer graphs).
@@
Next, we proceed to compare the performance of GVE-LPA with GVE-Louvain (our parallel implementation of Louvain algorithm) in Figure 5. Figures 5(a), 5(b), and 5(c) present the runtimes, speedup (of GVE-LPA with respect to GVE-Louvain), and modularity of GVE-Louvain and GVE-LPA for each graph in the dataset. GVE-LPA offers a mean speedup of 5.4×5.4$\times$5.4 × compared to GVE-Louvain, particularly on social networks, road networks, and protein k-mer graphs, while achieving on average $10.9\%$ lower modularity, especially on social networks and protein k-mer graphs.
In Figure 6, we assess the runtime per edge of GVE-LPA. Results indicate that GVE-LPA exhibits a higher $\text{runtime}/|E|$ factor on graphs with a low average degree (road networks and protein k-mer graphs) and on graphs with a weakly connected clusters (com-LiveJournal and com-Orkut graphs).
@@ 6. Conclusion
In summary, this study focuses on optimizing the Label Propagation Algorithm (LPA), a high-speed community detection algorithm, in the shared memory setting. We consider 6 different optimizations, which significantly improve the performance of the algorithm. Comparative assessments against competitive open-source implementations (FLPA) and packages (igraph and NetworKit) indicate that GVE-LPA outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97000×97000$\times$97000 ×, and 40×40$\times$40 × respectively, while identifying communities of $6.6\%$ / $0.2\%$ higher quality than FLPA and igraph LPA respectively, and of $4.1\%$ lower quality than NetworKit LPA.
On a web graph with $3.8$ billion edges, GVE-LPA identifies communities in $2.7$ seconds, and thus achieves a processing rate of $1.4$ billion edges/s. GVE-LPA is thus ideal for applications requiring a high-speed clustering algorithm while accepting a compromise in clustering quality — it is on average 5.4×5.4$\times$5.4 × faster than GVE-Louvain, and has a strong scaling factor of 1.7×1.7$\times$1.7 × for every doubling of threads, but identifies communities with on average $10.9\%$ lower modularity than GVE-Louvain.